# PromptBench: A Unified Library for Evaluation of Large Language Models

### 13 Dec 2023 v1

@@ Abstract
The evaluation of large language models (LLMs) is crucial to assess their performance and mitigate potential security risks. In this paper, we introduce PromptBench, a unified library to evaluate LLMs. It consists of several key components that are easily used and extended by researchers: prompt construction, prompt engineering, dataset and model loading, adversarial prompt attack, dynamic evaluation protocols, and analysis tools. PromptBench is designed to be an open, general, and flexible codebase for research purpose that can facilitate original study in creating new benchmarks, deploying downstream applications, and designing new evaluation protocols. The code is available at: https://github.com/microsoft/promptbench and will be continuously supported.
@@ 1 Introduction
Large language models (LLMs) have been revolutionizing numerous applications, impacting virtually every aspect of human life and society unprecedentedly. They are trained on a massive scale of carefully curated corpora (Brown et al., 2020; Team, 2023; Computer, 2023), enabling humans to interact with them easily and flexibly through natural language to solve complicated real-world tasks (OpenAI, 2023b) through prompts.
@@
Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023). Recent efforts have evaluated LLMs from diverse aspects (Liang et al., 2022; Zheng et al., 2023; Li et al., 2023b; Huang et al., 2023). Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022). Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation.
@@
Existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022) offer an entire LLM pipeline for various downstream applications, they are not built inherently for evaluation purpose. LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling advanced, context-aware functionalities. Semantic Kernel aims to merge AI services with programming languages for versatile AI app development. The eval-harness (Gao et al., 2023) library offers a comprehensive framework for evaluating generative language models, but it does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation. Consequently, there is an urgent need for a unified library dedicated to the comprehensive evaluation of diverse aspects, particularly for research purposes.
@@
This paper introduces PromptBench, a unified python library for evaluating LLMs from comprehensive dimensions. [^1] The name “PromptBench” is the same as (Zhu et al., 2023b) where we only evaluated the robustness against adversarial prompt attack. We decided to keep the name and heavily extend that project. It consists of a wide range of LLMs and evaluation datasets, covering diverse tasks, evaluation protocols, adversarial prompt attacks, and prompt engineering techniques. As a holistic library, it also supports several analysis tools for interpreting the results. Our library is designed in a modular fashion, allowing researchers to easily build evaluation pipelines for their own projects. We open-source PromptBench with comprehensive documents and tutorials [^2] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation. We believe PromptBench could boost our understanding of capabilities of LLMs and facilitate new research from the community.
@@ 2.1 Components
Models. PromptBench supports both open-source and proprietary LLMs and it is open to add more. Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b). PromptBench provides a unified LLMModel interface to allow easy construction and inference of a model with specified max generating tokens and generating temperature. More details of the supported models are shown in Appendix A.1.
@@ 2.2 Evaluation pipeline
PromptBench allows easy construction of an evaluation pipeline via four steps. Firstly, specify task and then load dataset via pb.DatasetLoader. PromptBench offers a streamlined one-line API for loading the desired dataset. Secondly, users can customize LLMs using the pb.LLMModel, which provides integrated inference pipelines compatible with most LLMs implemented in Huggingface. Thirdly, the prompt for the specified dataset and task is defined via pb.Prompt. Users have the option to input a list of prompts for evaluation and performance comparison. In cases where no prompts are supplied, our default prompts for the dataset are utilized. Finally, the pipeline requires the definition of input and output processing functions via class InputProcess and class OutputProcess defined in pb.utils.dataprocess, as well as the evaluation function via pb.metrics. The detailed introduction of the components are shown in Appendix A.
@@ 2.3 Supported research topics
PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix B. Researchers are welcome to submit new results to our platform. Extensibility is shown in Appendix C that allows convenient extension of the framework.
@@ 3 Conclusion
This work presented PromptBench, a unified framework for evaluation of large language models. The library is designed in modular fashion to allow users build evaluation pipeline by composing different models, tasks, and prompts. It also facilitates several research directions such as prompt engineering, adversarial prompt attacks, and dynamic evaluation. We treat PromptBench as the first step towards assessing the true capabilities and exploring the boundaries of current LLMs, and believe the benchmark and analysis results from our library could shed lights on designing more robust and human-aligned models. PromptBench is a long-term project and will be actively updated to fulfill the needs from the research community. We welcome contributions from all potential contributors.
Appendix A Details of PromptBench
A.1 Models
@@
  • phi-1.5 (Li et al., 2023c): phi-1.5 is a transformer architecture with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
@@
  • PaLM 2 (Anil et al., 2023) PaLM 2 is an advanced language model that excels in multilingual and reasoning capabilities, offering greater computational efficiency than its predecessor, PaLM. This Transformer-based model enhances performance across various model sizes in English, multilingual tasks, and reasoning challenges.
A.2 Tasks and Datasets
@@
  • GLUE (Wang et al., 2019) The GLUE benchmark (General Language Understanding Evaluation) offers a suite of tasks to evaluate the capability of NLP models in understanding language. For this research, we employed 8 specific tasks: Sentiment Analysis (SST-2 (Socher et al., 2013)), Grammar Correctness (CoLA (Warstadt et al., 2018)), Identifying Duplicate Sentences (QQP (Wang et al., 2017), MRPC (Dolan and Brockett, 2005)), and various Natural Language Inference tasks (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), WNLI (Levesque et al., 2012)).
@@
  • MMLU (Hendrycks et al., 2021) The MMLU dataset tests the broad knowledge and problem-solving skills of large language models through 57 tasks with multiple-choice questions in fields like mathematics, history, and computer science. It is a comprehensive multitask benchmark.
@@
  • SQuAD V2 (Rajpurkar et al., 2018) The SQuAD v2 dataset is pivotal in training and assessing NLP models for reading comprehension. It builds upon the original SQuAD by adding unanswerable questions, making it more challenging. Models must either identify the correct answer in the text or recognize questions as unanswerable.
@@
  • UN Multi (Eisele and Chen, 2010) Comprising texts in the six official United Nations languages, the Multi UN dataset is a vast parallel corpus from UN documents. However, its focus on formal texts may restrict its use in informal or conversational language contexts.
@@
  • IWSLT 2017 (Cettolo et al., 2017) Designed for spoken language translation system evaluation, the IWSLT 2017 dataset includes multilingual, multi-domain text data, primarily from the TED Talks Open Translation Project. It encompasses numerous language pairs, providing a rich resource for translation tasks.
@@
  • Math (Saxton et al., 2019) The DeepMind Mathematics Dataset assesses AI models’ mathematical reasoning by posing a wide array of math problems, from algebra to calculus. It tests the models’ understanding and logical reasoning in mathematics.
@@
  • BigBench (bench authors, 2023) The BIG-bench is a collaborative benchmark designed to evaluate the capabilities of large language models and predict their future potential. It consists of over 200 tasks, contributed by 444 authors from 132 institutions, covering a wide range of topics like linguistics, math, common-sense reasoning, and more. These tasks are intended to probe areas believed to be beyond the current capabilities of language models.
@@
  • GSM8K (Cobbe et al., 2021) The GSM8K dataset is a collection of 8.5K high-quality, linguistically diverse grade school math word problems. It was created by human problem writers and is divided into 7.5K training problems and 1K test problems. These problems, which require 2 to 8 steps to solve, primarily involve basic arithmetic operations and are designed to be solvable by a bright middle school student.
@@
  • CommonsenseQA (Talmor et al., 2019) The CommonsenseQA dataset is a challenging commonsense question-answering dataset. It comprises 12,247 questions with 5 multiple-choice answers each.
@@
  • QASC (Khot et al., 2020) QASC (Question Answering via Sentence Composition) is a specialized collection designed for question-answering tasks with a focus on sentence composition. It comprises 9,980 eight-way multiple-choice questions about grade school science, divided into 8,134 for training, 926 for development, and 920 for testing .(In our evaluation, we use development part.) The dataset is notable for its emphasis on multi-hop reasoning, requiring the retrieval and composition of facts from a broad corpus to answer each question.
@@
  • NummerSense (Lin et al., 2020) NumerSense is a unique numerical commonsense reasoning probing task, featuring a diagnostic dataset with 3,145 masked-word-prediction probes. This dataset has applications in tasks such as knowledge base completion and open-domain question answering.
A.3 Evaluation protocols
@@ A.4.2 Prompt Engineering
A.4.3 Adversarial Prompt Attacks
@@ A.5 Pipeline
Appendix B Benchmark Results
B.1 Adversarial prompt robustness
@@ C.2 Add new models
    Implementing a New Model Class: Models should be implemented in dataload/model.py, inheriting from the LLMModel class. In your customized model, you should implement self.tokenizer and self.model. You may also customize your own predict function for inference. If the predict function is not customized, the default predict function inherited from LLMModel will be used.
@@
    Adding an Interface: After customizing the model class, register it in the _create_model function within the class LLMModel in __init__.py.
C.3 Add new prompt engineering methods
@@
C.4 Add new metrics and input/output process functions

### 05 Jan 2024 v2

@@ Abstract
The evaluation of large language models (LLMs) is crucial to assess their performance and mitigate potential security risks. In this paper, we introduce PromptBench, a unified library to evaluate LLMs. It consists of several key components that can be easily used and extended by researchers: prompt construction, prompt engineering, dataset and model loading, adversarial prompt attack, dynamic evaluation protocols, and analysis tools. PromptBench is designed as an open, general, and flexible codebase for research purpose. It aims to facilitate original study in creating new benchmarks, deploying downstream applications, and designing new evaluation protocols. The code is available at: https://github.com/microsoft/promptbench and will be continuously supported.
@@ 1 Introduction
Large language models (LLMs) have been revolutionizing various aspects of human life and society unprecedentedly. Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023). Recent efforts have evaluated LLMs from diverse aspects (Liang et al., 2022; Zheng et al., 2023; Li et al., 2023b; Huang et al., 2023). Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022). On top of that, there have been various prompt learning algorithms developed based on different evaluation metrics, such as BDPL (Diao et al., 2022), GrIPS (Prasad et al., 2022) and Plum (Pan et al., 2023). Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation.
@@
There are existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022). However, LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling context-aware functionalities. Semantic Kernel aims to merge AI services with programming languages for versatile AI app development. Eval-harness (Gao et al., 2023) offers a comprehensive framework for evaluating generative language models, but does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation. Zeno (Zeno, 2023) is an AI evaluation platform supporting interaction and visualization, but it is not easy to customize. LiteLLM (BerriAI, 2023) implements a unified API call for different LLM service prodiders, but it does not support research topics such as robustness, prompt engineering, and new protocols. Detailed comparisons are shown in Appendix A. Consequently, there is an urgent need for a unified library dedicated to the comprehensive evaluation of diverse aspects, particularly for research purposes.
@@
This paper introduces PromptBench, a unified python library for evaluating LLMs from comprehensive dimensions. [^1] The name “PromptBench” is the same as (Zhu et al., 2023b) where we only evaluated the robustness against adversarial prompt attack. We decided to keep the name and heavily extend that project. It consists of a wide range of LLMs and evaluation datasets, covering diverse tasks, evaluation protocols, adversarial prompt attacks, and prompt engineering techniques. As a holistic library, it also supports several analysis tools for interpreting the results. Our library is designed in a modular fashion, allowing researchers to easily build evaluation pipelines for their own projects. We open-source PromptBench with comprehensive documents and tutorials [^2] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation. We believe PromptBench could enhance our understanding of LLMs’ capabilities and spur new research within the community.
@@ 2.1 Components
Models. PromptBench supports both open-source and proprietary LLMs and it is open to add more. Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 series (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b). PromptBench provides a unified LLMModel interface to allow easy construction and inference of a model with specified max generating tokens and generating temperature. More details of the supported models are shown in Appendix B.1.
@@ 2.2 Evaluation pipeline
PromptBench allows easy construction of an evaluation pipeline via four steps. Firstly, specify task and then load dataset via pb.DatasetLoader. PromptBench offers a streamlined one-line API for loading the desired dataset. Secondly, users can customize LLMs using the pb.LLMModel, which provides integrated inference pipelines compatible with most LLMs implemented in Huggingface. Thirdly, the prompt for the specified dataset and task is defined via pb.Prompt. Users have the option to input a list of prompts for evaluation and performance comparison. In cases where no prompts are supplied, our default prompts for the dataset are utilized. Finally, the pipeline requires the definition of input and output processing functions via class InputProcess and class OutputProcess defined in pb.utils.dataprocess, as well as the evaluation function via pb.metrics. The detailed introduction of the components are shown in Appendix B.
@@ 2.3 Supported research topics
PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix C. Researchers are welcome to submit new results to our platform. Extensibility is shown in Appendix D that allows convenient extension of the framework.
@@ 3 Conclusion
We presented PromptBench, a unified framework for LLMs evaluation. The library is designed in modular fashion to allow users build evaluation pipeline by composing different models, tasks, and prompts. It also facilitates several research directions such as prompt engineering, adversarial prompt attacks, and dynamic evaluation. We treat PromptBench as the first step towards assessing the true capabilities and exploring the boundaries of current LLMs, and believe the benchmark and analysis results from our library could shed lights on designing more robust and human-aligned models. PromptBench is a long-term project and will be actively maintained. We welcome contributions from all potential contributors.
Appendix A Comparison with Related Code Libraries
Table 1: Comparison with related code libraries.
Library Purpose Customization Functions
OpenAI Evals (Contributors, 2023b)
A framework for evaluating LLMs or systems built using LLMs.
model, dataset, prompt, eval
Evaluation pipelines Benchmarks
OpenCompass (Contributors, 2023c)
One-stop platform for large model evaluation, aiming to provide a transparent benchmark for foundation models.
model, dataset
Evaluation pipelines Benchmarks Leaderboards
promptfoo (Contributors, 2023a)
A framework for evaluating prompts and large language models.
model, dataset, prompt, eval
Evaluation pipelines
LM Evaluation Harness (Gao et al., 2023)
A framework for evaluation of autoregressive LMs.
model, dataset, prompt, eval
Evaluation pipelines Benchmarks Leaderboards
HELM (Liang et al., 2022)
Holistic Evaluation of Language Models.
model
Evaluation pipelines Benchmark Leaderboard
PromptBench (Ours)
Research-focused evaluation toolkit.
model, dataset, prompt, prompt engineering, eval
Evaluation pipelines Prompt Engineering Prompt attacks Dynamic evaluation Leaderboard
Appendix B Details of PromptBench
B.1 Models
@@
  • phi-1.5 and phi-2 (Li et al., 2023c): phi-1.5 is an LLM with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
  • Mistral 7B (Jiang et al., 2023): Mistral 7B is trained by Mistral AI team. It excels in tasks like reasoning, mathematics, and code generation. It uses grouped-query attention for faster inference and sliding window attention for efficient handling of sequences. There’s also an instruction-following version, Mistral 7B-Instruct.
  • Mixtral8x7B (Mixtral, 2023): Engineering by Mistral AI team, this model is a high-quality sparse mixture of experts model (SMoE) with open weights. There’s also an instruction-following version, Mixtral 8x7B Instruct.
  • Baichuan2 series (Yang et al., 2023): Baichuan2 is developed by Baichuan Intelligent. Trained on 2.6 trillion high-quality tokens, it achieves the best results in its size class on multiple authoritative benchmarks in Chinese, English, and multilingual general and domain-specific tasks.
  • Yi series (Yi, 2023): Developed by 01.AI, the Yi series are next-generation open-source large language models. Trained on a 3T multilingual corpus, they excel in language understanding, commonsense reasoning, and reading comprehension.
Proprietary LLMs:
@@
  • PaLM 2 (Anil et al., 2023): PaLM 2 is an advanced language model that excels in multilingual and reasoning capabilities, offering greater computational efficiency than its predecessor, PaLM. This Transformer-based model enhances performance across various model sizes in English, multilingual tasks, and reasoning challenges.
  • Gemini (Google, 2023): Gemini is Google’s new family of LLMs consisting of Gemini Nano, Gemini Pro, and Gemini Ultra. Gemini Pro is released via API service.
B.2 Tasks and Datasets
@@
  • GLUE (Wang et al., 2019): The GLUE benchmark (General Language Understanding Evaluation) offers a suite of tasks to evaluate the capability of NLP models in understanding language. For this research, we employed 8 specific tasks: Sentiment Analysis (SST-2 (Socher et al., 2013)), Grammar Correctness (CoLA (Warstadt et al., 2018)), Identifying Duplicate Sentences (QQP (Wang et al., 2017), MRPC (Dolan and Brockett, 2005)), and various Natural Language Inference tasks (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), WNLI (Levesque et al., 2012)).
@@
  • MMLU (Hendrycks et al., 2021): The MMLU dataset tests the broad knowledge and problem-solving skills of large language models through 57 tasks with multiple-choice questions in fields like mathematics, history, and computer science. It is a comprehensive multitask benchmark.
@@
  • SQuAD V2 (Rajpurkar et al., 2018): The SQuAD v2 dataset is pivotal in training and assessing NLP models for reading comprehension. It builds upon the original SQuAD by adding unanswerable questions, making it more challenging. Models must either identify the correct answer in the text or recognize questions as unanswerable.
@@
  • UN Multi (Eisele and Chen, 2010): Comprising texts in the six official United Nations languages, the Multi UN dataset is a vast parallel corpus from UN documents. However, its focus on formal texts may restrict its use in informal or conversational language contexts.
@@
  • IWSLT 2017 (Cettolo et al., 2017): Designed for spoken language translation system evaluation, the IWSLT 2017 dataset includes multilingual, multi-domain text data, primarily from the TED Talks Open Translation Project. It encompasses numerous language pairs, providing a rich resource for translation tasks.
@@
  • Math (Saxton et al., 2019): The DeepMind Mathematics Dataset assesses AI models’ mathematical reasoning by posing a wide array of math problems, from algebra to calculus. It tests the models’ understanding and logical reasoning in mathematics.
@@
  • BIG-Bench (bench authors, 2023): BIG-bench is a collaborative benchmark designed to evaluate the capabilities of large language models and predict their future potential. It consists of over 200 tasks, contributed by 444 authors from 132 institutions, covering a wide range of topics like linguistics, math, common-sense reasoning, and more. These tasks are intended to probe areas believed to be beyond the current capabilities of LMs.
@@
  • GSM8K (Cobbe et al., 2021): The GSM8K dataset is a collection of 8.5K high-quality, linguistically diverse grade school math word problems. It was created by human problem writers and is divided into 7.5K training problems and 1K test problems. These problems, which require 2 to 8 steps to solve, primarily involve basic arithmetic operations and are designed to be solvable by a bright middle school student.
@@
  • CommonsenseQA (Talmor et al., 2019): The CommonsenseQA dataset is a challenging commonsense question-answering dataset. It comprises 12,247 questions with 5 multiple-choice answers each.
@@
  • QASC (Khot et al., 2020): QASC (Question Answering via Sentence Composition) is a specialized collection designed for question-answering tasks with a focus on sentence composition. It comprises 9,980 eight-way multiple-choice questions about grade school science, divided into 8,134 for training, 926 for development, and 920 for testing .(In our evaluation, we use development part.) The dataset is notable for its emphasis on multi-hop reasoning, requiring the retrieval and composition of facts from a broad corpus to answer each question.
@@
  • NummerSense (Lin et al., 2020): NumerSense is a unique numerical commonsense reasoning probing task, featuring a diagnostic dataset with 3,145 masked-word-prediction probes. This dataset has applications in tasks such as knowledge base completion and open-domain question answering.
B.3 Evaluation protocols
@@ B.4.2 Prompt Engineering
Note that there are plenty of prompt engineering techniques and we tried our best to include those general techniques instead of specific prompt engineering techniques such as Tree of Thoughts (Yao et al., 2023) that requires specific prompt design and decomposition of each problem.
B.4.3 Adversarial Prompt Attacks
@@ B.5 Pipeline
Appendix C Benchmark Results
C.1 Adversarial prompt robustness
@@ D.2 Add new models
    Implementing a New Model Class: Models should be implemented in models/model.py, inheriting from the LLMModel class. In your customized model, you should implement self.tokenizer and self.model. You may also customize your own predict function for inference. If the predict function is not customized, the default predict function inherited from LLMModel will be used.
@@
    Adding an Interface: After customizing the model class, register it in the _create_model function within the class LLMModel and MODEL_LIST dictionary in __init__.py.
D.3 Add new prompt engineering methods
@@
D.4 Add new metrics and input/output process functions